OpenStudy (anonymous):

Algebra 2. *******MEDAL*******

OpenStudy (anonymous):

7x + 2y = 4 y = x + 1 Solve the system of equations. A. (1/3, 4/3) B. (2/9, 11/9) C. no solution

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Can you tell me how you got that?

OpenStudy (anonymous):

7x+2y=4 7(2/9)+2(11/9)=4 1+(5/9)+2+(4/9)=4 3+(9/9)=4 4=4 y=x+1 (11/9)=(2/9)+1 (11/9)=(11/9)